OpenStudy (anonymous):

The highest temperature recorded in a city last year was ninety-two degrees Fahrenheit . The lowest temperature recorded in the same city last year negative five degrees Fahrenheit . How much greater was last year's record high temperature than the record low for this city?

OpenStudy (yttrium):

It's easy! :)) Just subtract the lowest recorded temp to the highest recorded temp. Mathematically: 92-(-5)= ???

OpenStudy (anonymous):

just the difference of the two i:e ( 92- (-5)) =(92+5) ...

OpenStudy (anonymous):

so its 97

OpenStudy (yttrium):

Gotcha.

OpenStudy (anonymous):

right

OpenStudy (anonymous):

thanks but idk who to givee the best response to lol

OpenStudy (anonymous):

its ok..good luck